OpenStudy (anonymous):

2.42/7.86 = 3.59/ ? what steps should i follow to solve

hero (hero):

\[\left(2 + \frac{42}{100}\right) \div \left(7 + \frac{86}{100}\right)\]\[\left(\frac{200}{100} + \frac{42}{100}\right) \div \left(\frac{700}{100} + \frac{86}{100}\right)\]\[\left(\frac{242}{100}\right) \div \left(\frac{786}{100}\right)\]\[\left(\frac{242}{100}\right) \times \left(\frac{100}{786}\right)\]

hero (hero):

The 100's cancel and you're just left with \[\frac{242}{786}\] You can reduce that fraction to lowest terms.

hero (hero):

Oh wait, did you mean \[\frac{2.42}{7.86} = \frac{3.59}{x}\] And you have to solve for x?

hero (hero):

If that's what you meant, then you should cross multiply, and then isolate x. 2.42x = (3.59)(7.86)

OpenStudy (anonymous):

good! i was lost ... il try that

hero (hero):

Okay, good luck with it.

OpenStudy (anonymous):

i do nothing with 2.42 correct ?

hero (hero):

You divide both sides by the 2.42

hero (hero):

The goal is to isolate x...in other words to get x by itself.

hero (hero):

x is not isolated with the 2.42 attached to it.

hero (hero):